I am confused with this problem. Can someone help?
f(x)=5x-1
Find the followinf for f(a-2.
Thanks.

Answer by ankor@dixie-net.com(22740)

f(x)=5x-1
Find the following for f(a-2)
:
Just substitute (a-2) for x in the original equation.
f(a-2) = 5(a-2) - 1
Multiply what's inside the brackets:
f(a-2( = 5a - 10 - 1
:
f(a-2) = 5a - 11